suppose a uniform electric field of magnitude 150 n/c exists in a region of space. a) how far apart are a pair of equipotential surfaces whose potentials differ by 6.0 V? b) draw a sketch of the electric field lines and the equipotential sufaces
Answer #1

a) Electric field = Voltage / Difference

E = V/d

d = V/E = 6 / 150 = 0.04 m

= 4cm

so the seperation is 4 cm